OpenStudy (anonymous):

Calculate the steepness of the line passing through the points (4, 24) and (10, 12). 3 -2 2 2.5

OpenStudy (anonymous):

by steepness do you mean slope?

OpenStudy (anonymous):

i guess so

OpenStudy (princeharryyy):

12-24 / 10 -4 = -12/6 = -2